The time lag between 2 particles vibrating in a progressive wave separated by a distance 20m is 0.02s . The wave velocity of he frequency of the wave is 500hz,is
 

Grade:12

2 Answers

Nallamilli Veena Madhuri
6 years ago
Velocity=frequency×wavelength=500× (1/500×0.02)×20=20/0.02=1000m/s , how to find length,for 0.02sec---20m and for T=1/f=1/500------? ?=1/(500×0.02) ×20m= wavelength
Karthik
5 years ago
 Here , Velocity = distance covered / time taken
 c=20/0.02=1000ms-1 . So , wave velocity is 1000ms-1 . This can be bone in other way too by converting the Hz value